# Smooth Lower Bounds for Differentially Private Algorithms via Padding-and-Permuting Fingerprinting Codes

### v3

@@ 7.2 $k$-Means Clustering
We say that $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k}$ is an $(\lambda,\xi,\beta)$-approximation algorithm for $k$-means, if for every $S\in({\cal B}_{d})^{n}$ it holds that
${\mathrm{Pr}}_{C\sim\mathsf{M}(S)}\mathopen{}\mathclose{{}\left[{\rm COST}_{S}(C)\leq\lambda\cdot{\rm OPT}_{k}(S)+\xi}\right]\geq\beta$
Theorem 7.2 (Our Lower Bound, Restatement of Theorem 1.4).
Let $n,d,k\in{\mathbb{N}}$, $\lambda\geq 1$ and $\xi\geq 0$ such that $k\geq 2$ and $n\geq k+\xi$. If $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k}$ is an $(1,\frac{\beta}{4nk})$-DP $(\lambda,\xi,\beta)$-approximation algorithm for $k$-means, then either $k\geq 2^{\Omega(d/\lambda)}\beta\lambda/d$ or $\xi\geq\Omega\mathopen{}\mathclose{{}\left(\frac{k\sqrt{d/\lambda}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{kd}{\beta\lambda}}\right)}}\right)$.
@@
Let $n,d,k\in{\mathbb{N}}$, $\lambda\geq 1$ and $\xi\geq 1$ such that $n\geq m$ for $m=\mathopen{}\mathclose{{}\left\lfloor 1+\xi/k}\right\rfloor\cdot k$. If $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k+1}$ is an $(\lambda,\>\xi,\>\beta)$-approximation algorithm for $(k+1)$-means, then the mechanism $\widetilde{\mathsf{M}}\colon(\{-1,1\}^{d})^{m}\rightarrow\{-1,1\}^{d}$ defined in Section 7.2 is $(k,\>\alpha=\frac{1}{40\lambda},\>\frac{\beta}{k+1})$-weakly-accurate (Section 1.1.1).
@@
Input: $x_{1},\ldots,x_{m}\in\{-1,1\}^{d}$ for $m=\mathopen{}\mathclose{{}\left\lfloor 1+\xi/k}\right\rfloor\cdot k$.
Operation:
  1. 1.
    Compute $(c_{1},\ldots,c_{k+1})=\mathsf{M}(\frac{1}{\sqrt{d}}x_{1},\ldots,\frac{1}{\sqrt{d}}x_{m},\underbrace{\vec{0},\ldots,\vec{0}}_{n-m\text{ times}})$ (where $\vec{0}=(\underbrace{0,\ldots,0}_{d\text{ times}})$).
  2. 2.
    Sample $j\leftarrow[k+1]$ and output $\mathsf{sign}\mathopen{}\mathclose{{}\left(c_{j}}\right)$.
We first prove Theorem 7.2 using Section 7.2.
@@
Let $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k}$ be the mechanism from Theorem 7.2. By Section 7.2, the mechanism $\widetilde{\mathsf{M}}\colon(\{-1,1\}^{d})^{m}\rightarrow\{-1,1\}^{d}$ (Section 7.2) is $(k-1,\alpha=\frac{1}{40\lambda},\frac{\beta}{k})$-weakly-accurate. Furthermore, by post-processing, $\widetilde{\mathsf{M}}$ is also $\mathopen{}\mathclose{{}\left(1,\frac{\beta}{4nk}}\right)$-DP, which in particular implies that it is $\mathopen{}\mathclose{{}\left(1,\frac{\beta}{4n(k-1)}}\right)$-DP. Hence we deduce by Theorem 1.2 that $m\geq\Omega\mathopen{}\mathclose{{}\left(\frac{k\sqrt{d/\lambda}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{dk}{\beta\lambda}}\right)}}\right)$. Since $m\geq\max\{k,\xi\}$, the above implies that either $k\geq\Omega\mathopen{}\mathclose{{}\left(\frac{k\sqrt{d/\lambda}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{dk}{\beta\lambda}}\right)}}\right)$ (which is translated to $k\geq 2^{\Omega(d/\lambda)}\beta\lambda/d$), or that $\xi\geq\Omega\mathopen{}\mathclose{{}\left(\frac{k\sqrt{d/\lambda}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{dk}{\beta\lambda}}\right)}}\right)$.     $\Box$
We now prove Section 7.2.
Proof of Section 7.2.
@@
$\mathopen{}\mathclose{{}\left\lVert c_{t}^{*}-\tilde{x}_{i}}\right\rVert_{2}^{2}=\frac{1}{d}\cdot\mathopen{}\mathclose{{}\left\lVert x_{tm/k}-x_{i}}\right\rVert_{2}^{2}\leq 4\alpha,$
@@
${\rm OPT}_{k+1}(S)\leq{\rm COST}_{S}(c_{1}^{*},\ldots,c_{k+1}^{*})\leq 4\alpha m.$
Since $\lambda\cdot{\rm OPT}_{k+1}(S)+\xi\leq m/40$, the utility guarantee of $\mathsf{M}$ implies that
${\mathrm{Pr}}_{C=(c_{1},\ldots,c_{k+1})\sim\mathsf{M}(S)}\mathopen{}\mathclose{{}\left[{\rm COST}_{S}(C)\leq m/40}\right]\geq\beta.$
We prove the claim by showing that for every $C=(c_{1},\ldots,c_{k+1})$ with ${\rm COST}_{S}(C)\leq m/40$, there exists $s\in[k+1]$ and $t\in[k]$ such that $\mathsf{sign}(c_{s})$ strongly-agrees with $X_{t}$. This will conclude the proof since the probability that the random $j$ chosen in Step 2 of $\widetilde{\mathsf{M}}$ will hit the right $s$ is $1/(k+1)$. Indeed if this is not the case, then for any center $c_{s}$ and any non-zero point $\tilde{x}_{i}\in{\cal S}$ where $x_{i}\in X_{t}$ we have at least one $b\in\{-1,1\}$ such that |$\mathopen{}\mathclose{{}\left|\{j\in{\cal J}_{X_{t}}^{b}\colon\mathsf{sign}(\tilde{x}_{i}^{j})\neq\mathsf{sign}(c_{s}^{j})\}}\right|>0.1\cdot\mathopen{}\mathclose{{}\left|{\cal J}_{X_{t}}^{b}}\right|>\frac{1}{20}(1-\alpha)d$, which yields that
$\mathopen{}\mathclose{{}\left\lVert\tilde{x}_{i}-c_{s}}\right\rVert_{2}^{2}\geq\frac{1}{d}\cdot\mathopen{}\mathclose{{}\left|\{j\in[d]\colon\mathsf{sign}(\tilde{x}_{i}^{j})\neq\mathsf{sign}(c_{s}^{j})\}}\right|\geq\frac{1}{d}\cdot\mathopen{}\mathclose{{}\left|\{j\in{\cal J}_{X_{t}}^{b}\colon\mathsf{sign}(\tilde{x}_{i}^{j})\neq\mathsf{sign}(c_{s}^{j})\}}\right|>\frac{1}{20}(1-\alpha).$
Thus ${\rm COST}_{C}(S)>\frac{m}{20}(1-\alpha)>m/40$, a contradiction to the assumption ${\rm COST}_{S}(\tilde{C})\leq m/40$. This concludes the proof.
@@ 7.3 Top Singular Vector
Fix $X=(x_{1},\ldots,x_{n})\in\{-1,1\}^{n\times d}$ with $\mathopen{}\mathclose{{}\left|{\cal J}^{1}_{X}}\right|,\mathopen{}\mathclose{{}\left|{\cal J}^{-1}_{X}}\right|\geq\frac{1}{2}(1-\alpha)d$, and let $\tilde{X}=\frac{1}{\sqrt{d}}X\in({\cal S}_{d})^{n}$. Let $u=(u_{1},\ldots,u_{d})\in{\cal S}_{d}$ be a column vector with $u_{j}=\begin{cases}1/\sqrt{d}&j\in{\cal J}^{1}_{X}\\ -1/\sqrt{d}&\text{o.w.}\end{cases}$. By definition it holds that $\mathopen{}\mathclose{{}\left\lVert\tilde{X}\cdot u}\right\rVert_{2}^{2}\geq\frac{n}{d}\mathopen{}\mathclose{{}\left(2\mathopen{}\mathclose{{}\left|{\cal J}_{X}^{1}}\right|+2\mathopen{}\mathclose{{}\left|{\cal J}_{X}^{-1}}\right|-d}\right)\geq(1-2\alpha)n$. This yields that $\sigma_{1}(\tilde{X})^{2}\geq(1-2\alpha)n$. Since $\sigma_{1}(\tilde{X})^{2}+\sigma_{2}(\tilde{X})^{2}\leq\mathopen{}\mathclose{{}\left\lVert\tilde{X}}\right\rVert_{F}^{2}=n$, we deduce that $\sigma_{2}(\tilde{X})^{2}/\sigma_{1}(\tilde{X})^{2}\leq\frac{2\alpha}{1-2\alpha}=\gamma^{2}$. Therefore, by the utility guarantee of $\mathsf{M}$, it holds that
@@
We prove the claim by showing that for any $y=(y_{1},\ldots,y_{d})\in{\cal S}_{d}$ with $\mathopen{}\mathclose{{}\left\lVert\tilde{X}\cdot y}\right\rVert_{2}^{2}\geq\mathopen{}\mathclose{{}\left\lVert\tilde{X}\cdot u}\right\rVert_{2}^{2}-\lambda\cdot\gamma n$, either $\mathsf{sign}(y)$ or $-\mathsf{sign}(y)$ strongly-agrees with $X$ (Section 1.1.1). In the following, fix such $y$, and let $s=\sum_{j\in{\cal J}_{X}^{1}}y_{j}-\sum_{j\in{\cal J}_{X}^{-1}}y_{j}$. We next prove that
@@
which is a contradiction to the assumption about $v$. We first prove Item 1. Let $z=(z_{1},\ldots,z_{n})=\tilde{X}\cdot y$ and let ${\cal B}=[d]\setminus({\cal J}_{X}^{1}\cup{\cal J}_{X}^{-1})$. Note that for every $i\in[n]$:
@@
where the last inequality holds since $\sum_{j\in{\cal B}}x_{i}^{j}\cdot y_{j}\leq\mathopen{}\mathclose{{}\left\lVert x_{i}^{{\cal B}}}\right\rVert_{2}\cdot\mathopen{}\mathclose{{}\left\lVert y_{{\cal B}}}\right\rVert_{2}\leq\sqrt{\mathopen{}\mathclose{{}\left|{\cal B}}\right|}\cdot 1\leq\sqrt{\alpha d}$. Using the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$ we deduce that $\mathopen{}\mathclose{{}\left\lVert z}\right\rVert_{2}^{2}=\sum_{i=1}^{n}z_{i}^{2}\leq\mathopen{}\mathclose{{}\left(\frac{2s^{2}}{d}+2\alpha}\right)n$. Hence
@@
where the third inequality holds since $\mathopen{}\mathclose{{}\left\lVert w}\right\rVert_{1}\leq\sqrt{m}\cdot\mathopen{}\mathclose{{}\left\lVert w}\right\rVert_{2}$ for every $w\in{\mathbb{R}}^{m}$, and the fourth one holds since $\mathopen{}\mathclose{{}\left\lVert y_{{\cal J}^{+}\cup{\cal J}_{X}^{-1}}}\right\rVert_{2}\leq\mathopen{}\mathclose{{}\left\lVert y}\right\rVert_{2}=1$. By the right inequality in (12), a similar calculation to Equation 13 yields that

### v4

@@ 7.2 Clustering
We say that $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k}$ is an $(\lambda,\xi,\beta)$-approximation algorithm for $(k,z)$-clustering, if for every $S\in({\cal B}_{d})^{n}$ it holds that
${\mathrm{Pr}}_{C\sim\mathsf{M}(S)}\mathopen{}\mathclose{{}\left[{\rm COST}_{z}(C;S)\leq\lambda\cdot{\rm OPT}_{k,z}(S)+\xi}\right]\geq\beta$
Theorem 7.2 (Our Lower Bound, Extension of Theorem 1.4).
Let $n,d,k\in{\mathbb{N}}$, $\lambda,z\geq 1$ and $\xi\geq 0$ such that $k\geq 2$ and $n\geq k+2\cdot 40^{z/2}\xi$. If $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k}$ is an $(1,\frac{\beta}{4nk})$-DP $(\lambda,\xi,\beta)$-approximation algorithm for $(k,z)$-clustering, then either $k\geq 2^{\Omega(d/\lambda^{2/z})}\beta\lambda^{2/z}/d\>$ or $\>\xi\geq\Omega\mathopen{}\mathclose{{}\left(\frac{2^{-O(z)}k\sqrt{d/\lambda^{2/z}}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{kd}{\beta\lambda^{2/z}}}\right)}}\right)$.
@@
Let $n,d,k\in{\mathbb{N}}$, $\lambda,z\geq 1$ and $\xi\geq 1$ such that $n\geq m$ for $m=k\cdot\mathopen{}\mathclose{{}\left\lfloor(1+40^{z/2}\cdot 2\xi/k)}\right\rfloor$. If $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k+1}$ is an $(\lambda,\>\xi,\>\beta)$-approximation algorithm for $(k+1,z)$-clustering, then the following mechanism $\widetilde{\mathsf{M}}\colon(\{-1,1\}^{d})^{m}\rightarrow\{-1,1\}^{d}$ is $(k,\>\alpha=\frac{1}{160\cdot(2\lambda)^{2/z}},\>\frac{\beta}{k+1})$-weakly-accurate (Definition 1.1.1).
@@
Input: $x_{1},\ldots,x_{m}\in\{-1,1\}^{d}$ for $m=\mathopen{}\mathclose{{}\left\lfloor(1+40^{z/2}\cdot 2\xi/k)}\right\rfloor\cdot k$. Operation: 1. Compute $(c_{1},\ldots,c_{k+1})=\mathsf{M}(\frac{1}{\sqrt{d}}x_{1},\ldots,\frac{1}{\sqrt{d}}x_{m},\underbrace{\vec{0},\ldots,\vec{0}}_{n-m\text{ times}})$ (where $\vec{0}=(\underbrace{0,\ldots,0}_{d\text{ times}})$). 2. Sample $j\leftarrow[k+1]$ and output $\mathsf{sign}\mathopen{}\mathclose{{}\left(c_{j}}\right)$.
We first prove Theorem 7.2 using Claim 7.2.
@@
Let $\mathsf{M}\colon({\cal B}_{d})^{n}\rightarrow({\cal B}_{d})^{k}$ be the mechanism from Theorem 7.2. By Claim 7.2, the mechanism $\widetilde{\mathsf{M}}\colon(\{-1,1\}^{d})^{m}\rightarrow\{-1,1\}^{d}$ is $(k-1,\alpha=\frac{1}{160\cdot(2\lambda)^{2/z}},\frac{\beta}{k})$-weakly-accurate. Furthermore, by post-processing, $\widetilde{\mathsf{M}}$ is also $\mathopen{}\mathclose{{}\left(1,\frac{\beta}{4nk}}\right)$-DP, which in particular implies that it is $\mathopen{}\mathclose{{}\left(1,\frac{\beta}{4n(k-1)}}\right)$-DP. Hence we deduce by Theorem 1.2 that $m\geq\Omega\mathopen{}\mathclose{{}\left(\frac{k\sqrt{d/\lambda^{2/z}}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{dk}{\beta\lambda^{2/z}}}\right)}}\right)$. Since $m\geq\max\{k,40^{z/2}\cdot 2\xi\}$, the above implies that either $k\geq\Omega\mathopen{}\mathclose{{}\left(\frac{k\sqrt{d/\lambda^{2/z}}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{dk}{\beta\lambda^{2/z}}}\right)}}\right)$ (which implies that $k\geq 2^{\Omega(d/\lambda^{2/z})}\beta\lambda^{2/z}/d\>$), or that $\xi\geq\Omega\mathopen{}\mathclose{{}\left(\frac{2^{-O(z)}k\sqrt{d/\lambda^{2/z}}}{\log^{1.5}\mathopen{}\mathclose{{}\left(\frac{dk}{\beta\lambda^{2/z}}}\right)}}\right)$.     $\Box$
We now prove Claim 7.2.
Proof of Claim 7.2.
@@
$\mathopen{}\mathclose{{}\left\lVert c_{t}^{*}-\tilde{x}_{i}}\right\rVert_{2}^{z}=\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left\lVert c_{t}^{*}-\tilde{x}_{i}}\right\rVert_{2}^{2}}\right)^{z/2}=\mathopen{}\mathclose{{}\left(\frac{1}{d}\cdot\mathopen{}\mathclose{{}\left\lVert x_{tm/k}-x_{i}}\right\rVert_{2}^{2}}\right)^{z/2}\leq(4\alpha)^{z/2},$
@@
${\rm OPT}_{k+1,z}(S)\leq{\rm COST}_{z}(c_{1}^{*},\ldots,c_{k+1}^{*};S)\leq(4\alpha)^{z/2}m.$
Now, since
$\lambda\cdot{\rm OPT}_{k+1,z}(S)+\xi$ $\leq\lambda\cdot(4\alpha)^{z/2}m+\frac{m}{2\cdot 40^{z/2}}$
$=\lambda\cdot\mathopen{}\mathclose{{}\left(\frac{1}{40(2\lambda)^{2/z}}}\right)^{z/2}+\frac{m}{2\cdot 40^{z/2}}$
$\leq m/40^{z/2},$
the utility guarantee of $\mathsf{M}$ implies that
${\mathrm{Pr}}_{C=(c_{1},\ldots,c_{k+1})\sim\mathsf{M}(S)}\mathopen{}\mathclose{{}\left[{\rm COST}_{z}(C;S)\leq m/40^{z/2}}\right]\geq\beta.$
We prove the claim by showing that for every $C=(c_{1},\ldots,c_{k+1})$ with ${\rm COST}_{z}(C;S)\leq m/40^{z/2}$, there exists $s\in[k+1]$ and $t\in[k]$ such that $\mathsf{sign}(c_{s})$ strongly-agrees with $X_{t}$. This will conclude the proof since the probability that the random $j$ chosen in Step 2 of $\widetilde{\mathsf{M}}$ will hit the right $s$ is $1/(k+1)$. Indeed if this is not the case, then for any center $c_{s}$ and any non-zero point $\tilde{x}_{i}\in S$ where $x_{i}\in X_{t}$ we have at least one $b\in\{-1,1\}$ such that |$\mathopen{}\mathclose{{}\left|\{j\in{\cal J}_{X_{t}}^{b}\colon\mathsf{sign}(\tilde{x}_{i}^{j})\neq\mathsf{sign}(c_{s}^{j})\}}\right|>0.1\cdot\mathopen{}\mathclose{{}\left|{\cal J}_{X_{t}}^{b}}\right|>\frac{1}{20}(1-\alpha)d$, which yields that
$\mathopen{}\mathclose{{}\left\lVert\tilde{x}_{i}-c_{s}}\right\rVert_{2}^{z}$ $=(\mathopen{}\mathclose{{}\left\lVert\tilde{x}_{i}-c_{s}}\right\rVert_{2}^{2})^{z/2}$
$\geq\mathopen{}\mathclose{{}\left(\frac{1}{d}\cdot\mathopen{}\mathclose{{}\left|\{j\in[d]\colon\mathsf{sign}(\tilde{x}_{i}^{j})\neq\mathsf{sign}(c_{s}^{j})\}}\right|}\right)^{z/2}$
$\geq\mathopen{}\mathclose{{}\left(\frac{1}{d}\cdot\mathopen{}\mathclose{{}\left|\{j\in{\cal J}_{X_{t}}^{b}\colon\mathsf{sign}(\tilde{x}_{i}^{j})\neq\mathsf{sign}(c_{s}^{j})\}}\right|}\right)^{z/2}$
$>\mathopen{}\mathclose{{}\left(\frac{1}{20}(1-\alpha)}\right)^{z/2}$
Thus ${\rm COST}_{z}(C;S)>\mathopen{}\mathclose{{}\left(\frac{1}{20}(1-\alpha)}\right)^{z/2}\cdot m>m/40^{z/2}$, a contradiction to the assumption ${\rm COST}_{z}(C;S)\leq m/40^{z/2}$. This concludes the proof of the claim.
@@ 7.3 Top Singular Vector
Fix $X=(x_{1},\ldots,x_{n})\in\{-1,1\}^{n\times d}$ with $\mathopen{}\mathclose{{}\left|{\cal J}^{1}_{X}}\right|,\mathopen{}\mathclose{{}\left|{\cal J}^{-1}_{X}}\right|\geq\frac{1}{2}(1-\alpha)d$, and let $\tilde{X}=\frac{1}{\sqrt{d}}X\in({\cal S}_{d})^{n}$. Let $u=(u_{1},\ldots,u_{d})\in{\cal S}_{d}$ be a column vector with $u_{j}=\begin{cases}1/\sqrt{d}&j\in{\cal J}^{1}_{X}\\ -1/\sqrt{d}&\text{o.w.}\end{cases}$. By definition it holds that $\mathopen{}\mathclose{{}\left\lVert\tilde{X}\cdot u}\right\rVert_{2}^{2}\geq\frac{n}{d}\mathopen{}\mathclose{{}\left(2\mathopen{}\mathclose{{}\left|{\cal J}_{X}^{1}}\right|+2\mathopen{}\mathclose{{}\left|{\cal J}_{X}^{-1}}\right|-d}\right)\geq(1-2\alpha)n$. This yields that $\sigma_{1}(\tilde{X})^{2}\geq(1-2\alpha)n$. Since $\sigma_{1}(\tilde{X})^{2}+\sigma_{2}(\tilde{X})^{2}\leq\mathopen{}\mathclose{{}\left\lVert\tilde{X}}\right\rVert_{F}^{2}=n$, we deduce that $\sigma_{2}(\tilde{X})^{2}/\sigma_{1}(\tilde{X})^{2}\leq\frac{2\alpha}{1-2\alpha}=\gamma^{2}$.
Therefore, by the utility guarantee of $\mathsf{M}$, it holds that
@@
We prove the claim by showing that for any $y=(y_{1},\ldots,y_{d})\in{\cal S}_{d}$ with $\mathopen{}\mathclose{{}\left\lVert\tilde{X}\cdot y}\right\rVert_{2}^{2}\geq\mathopen{}\mathclose{{}\left\lVert\tilde{X}\cdot u}\right\rVert_{2}^{2}-\lambda\cdot\gamma n$, either $\mathsf{sign}(y)$ or $-\mathsf{sign}(y)$ strongly-agrees with $X$ (Section 1.1.1).
In the following, fix such $y$, and let $s=\sum_{j\in{\cal J}_{X}^{1}}y_{j}-\sum_{j\in{\cal J}_{X}^{-1}}y_{j}$. We next prove that
@@
which is a contradiction to the assumption about $v$.
We first prove Item 1. Let $z=(z_{1},\ldots,z_{n})=\tilde{X}\cdot y$ and let ${\cal B}=[d]\setminus({\cal J}_{X}^{1}\cup{\cal J}_{X}^{-1})$. Note that for every $i\in[n]$:
@@
where the last inequality holds since $\sum_{j\in{\cal B}}x_{i}^{j}\cdot y_{j}\leq\mathopen{}\mathclose{{}\left\lVert x_{i}^{{\cal B}}}\right\rVert_{2}\cdot\mathopen{}\mathclose{{}\left\lVert y_{{\cal B}}}\right\rVert_{2}\leq\sqrt{\mathopen{}\mathclose{{}\left|{\cal B}}\right|}\cdot 1\leq\sqrt{\alpha d}$.
Using the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$ we deduce that $\mathopen{}\mathclose{{}\left\lVert z}\right\rVert_{2}^{2}=\sum_{i=1}^{n}z_{i}^{2}\leq\mathopen{}\mathclose{{}\left(\frac{2s^{2}}{d}+2\alpha}\right)n$. Hence
@@
where the third inequality holds since $\mathopen{}\mathclose{{}\left\lVert w}\right\rVert_{1}\leq\sqrt{m}\cdot\mathopen{}\mathclose{{}\left\lVert w}\right\rVert_{2}$ for every $w\in{\mathbb{R}}^{m}$, and the fourth one holds since $\mathopen{}\mathclose{{}\left\lVert y_{{\cal J}^{+}\cup{\cal J}_{X}^{-1}}}\right\rVert_{2}\leq\mathopen{}\mathclose{{}\left\lVert y}\right\rVert_{2}=1$.
By the right inequality in (12), a similar calculation to Equation 13 yields that